OpenStudy (anonymous):

Complete the square to form a perfect square trinomial x^2-12x+

OpenStudy (owlcoffee):

a perfect square is composed by: \[(a+b)^2= a^2 +ab + b^2\] in other words: "the first term squared, plus, two times the first times the second plus the second squared"

OpenStudy (anonymous):

I don't understand

OpenStudy (anonymous):

I don't understand

OpenStudy (owlcoffee):

All you have to do, is check the variables and see what you need to complete it. Since the structure was: \[(ax+b)^2=ax^2+2axb+b^2\] And you have: \[x^2-12x+...\] We can easily see that a=1, so we need a number, that makes 2ab=-12 So that be, since a=1: b=12/2 b=-6 So, now we have, in order to complete the perfect square, make the b^2, but since we deduced it was -6, we only square it: -6^2=36 so therefore: \[x^2-12x+36=(x-6)^2\]

OpenStudy (jhannybean):

Your function resembles the quadratic \(ax^2+bx+c\). You're looking for a C value to complete your quadratic function. To find it, use the formula: \[c=\left(\frac{b}{2}\right)^2\] In this case, \[c=\left(\color{red}{\frac{-12}{2}}\right)^2 =(\color{red}{-6})^2= 36\]To complete your quadratic, just plug in your c value. \[x^2-12x+36 \implies \left(x+\color{red}{\frac{b}{2}}\right)^2\implies (x\color{red}{-6})^2\]